# A genetic algorithm to generate maximally orthogonal frames in complex space

### v1

@@ I Introduction
As orthogonality is a pairwise property, increasing the orthogonality between a given pair of vectors will often come at the expense of decreasing the orthogonality between other pairs of vectors. Depending on how much value one assigns to the orthogonality of each pair, one will favor one “maximally orthogonal” set of vectors or another. Thus, the first step in this paper is providing a formal definition for maximally orthogonal frames of a complex space of arbitrary dimension. Once armed with this definition, we discuss their properties and establish their relation with notable frame families such as tight and Grassmannian frames, equiangular tight frames [10], symmetric, informationally complete, positive operator-valued measurements (SIC-POVMs) [11], maximal sets of mutually unbiased bases [12, 13] and complex projective $t$-designs [14, 15]. We also show that finding maximally orthogonal frames constitutes, in most cases, a numerical optimization problem. We discuss the relation with established numerical problems such as the Thomson problem [16], the problem of finding optimal packings in complex projetive space [17] and the problem of finding optimal reference states for quantum classifiers [18, 19]. Inspired by the use of a genetic algorithm for the Thomson problem [20], we present a genetic algorithm that is capable of generating maximally orthogonal frames of arbitrary size in a complex space of arbitrary dimension. Then, we visualize and measure the quality of the frames produced by the genetic algorithm. In the cases where the problem of finding maximally orthogonal frames is equivalent to the Thomson problem or the problem of finding optimal packings in complex projective space, we compare the genetic algorithm with state-of-the-art (SOTA) numerical methods.
@@ II.2.2 Riesz $s$-energy and asymptotic uniformity
The Riesz $s$-energy is invariant under changes in the global phase of the vectors, so we can identify equivalence classes for all unit vectors differing by a global phase, which essentially defines the complex projective space $\mathbb{C}P^{d-1}$. A parameterization of $\mathbb{C}P^{d-1}$ in real coordinates forms a manifold of dimension $2(d-1)$. Following the poppy-seed bagel theorem, we will have to set $s\geq 2(d-1)$ to generate MOFs, $\bar{\Phi}_{d,n}[s]$, that are asymptotically uniform, in the sense of uniformly distributed on the manifold that parametrizes $\mathbb{C}P^{d-1}$, for $n\to\infty$ [23].
@@ II.2.4 $1$-designs are tight
The frame potential was introduced by Bennedeto and Fickus, originally with $p=1$, for the study of finite normalized tight frames [28]. They proved that ${\rm FP}_{1}$ has several convenient properties: its global minima are tight frames and saturate the Welch bound ${\rm FP}_{1}(\bar{\Phi}_{d,n}[p=1])=n^{2}/d$, and all its local minima are degenerate, i.e. they are global minima and thus tight frames and $1$-designs. This tells us that we can see tight frames as a particular instance of MOFs as defined in Proposition II.1. In fact, Benedetto and Fickus already refer to a set of vectors that minimize the frame potential ${\rm FP}_{1}$ as maximally orthogonal. Thus, they equate MOFs with tight frames. As argued above, it is natural to relax this equation and consider tight frames as just one case of many MOFs. Just like ${\rm FP}_{1}$ leads to MOFs that are tight, other weighting functions are equally valid and endow the corresponding MOFs with other useful properties, such as uniformity or low coherence. We mention in passing that another consequence of Benedetto and Fickus’s work is that finding tight frames is an easy task. A local optimization from an arbitrary initial condition suffices to construct tight frames numerically, although constructive analytical methods have also been devised [29, 30]. Other weighting functions do not lead to such an easily minimizable energy function, and thus obtaining the corresponding MOFs poses a numerical challenge.
@@ II.2.5 Coherence: the $p\to\infty$ and $s\to\infty$ limits
$\lim_{s\to\infty}({\rm RE}_{s})^{1/s}=\alpha\max_{i\neq j}\frac{1}{D(|\phi_{i}\rangle,|\phi_{j}\rangle)}\,,$ (9)
where $\alpha^{s}$ is the multiplicity of the largest overlap. Therefore, only the maximum overlap(s) (minimum distance(s)) contribute to the energy and sets that only differ in submaximal overlaps will be degenerate. In this setting, the MOF will be given by
@@ II.2.6 Coherence and degeneracy
Figure 1: Degeneracy of the coherence. Bloch-sphere representation (top) and overlaps among vectors (bottom) of the different Grassmannian frames for $d=2$ and $n=5$. Column a) shows the distribution of vectors obtained by subtracting one vector from the equator from $\bar{\Phi}_{2,6}[s\to\infty]$. The frames depicted in columns b) and c) are obtained by continuously spreading the remaining three vectors at the equator until they form an equilateral triangle.
@@ II.2.7 Universal maximally orthogonal frames
The maximum number of MUB $\mathcal{M}(d)$ obeys $\mathcal{M}(d)\leq d+1$, with equality only if $d$ is an integer power of a primer number. Thus, if $d$ is an integer power of a prime it is possible to construct highly symmetric frames of $d(d+1)$ vectors, by combining the $d+1$ MUB, where every vector has only nearest-neighbors with overlap $1/d$ (the vectors of the other bases) and next-nearest-neighbors with null overlap (the other vectors of the same basis). Despite not being ETFs, frames built from $d+1$ sets of MUB are Grassmannian and tight [17] and complex projective $2$-designs [27].
@@ IV The genetic algorithm
Recombination is implemented as a random crossover of vectors from two individuals. A trivial recombination could consist on generating a random integer $i$ between $1$ and $n$ to create a child with the $i$ first vectors of parent A and the last $n-i$ vectors form parent B. However, the vectors within a frame are, a priori, not sorted in any particular fashion, i.e. their order does not reflect any geometrical structure in the manifold that parametrizes the vectors. Therefore, the subset of vectors inherited from each parent would be random, consisting of vectors from parent A that could be arbitrarily close to vectors from parent B, and thus throw away most of the orthogonality gained over the course of the optimization process. Because of this, a trivial recombination like the one just described would, most likely, create a child with a fitness much lower than that of its parents. To prevent this, we sort the vectors of each parent by distance to the first vector of each set (which is fixed and the same for all sets). After the sort, the order of the vectors in a set does reflect some structural properties. Namely, the first vector of the set marks a reference point in the manifold that is common for all sets. After sorting, we apply the random recombination by drawing the first $i$ vectors from parent A, which are expected to be relatively closer to the reference point, and the last $n-i$ vectors from parent B, which are expected to be relatively far from the reference point. To better understand this process, let us consider the case $d=2$. Now the manifold is the Bloch sphere and the reference point is given by the vector $(1,0)$, which marks the north pole. Selecting the first $i$ vectors from parent A corresponds to selecting vectors covering a spherical cap around the north pole with an approximate height determined by the ratio $i/n$. Then, selecting the $n-i$ vectors from parent B corresponds to selecting vectors covering a spherical cap around the south pole with an approximate height determined by the ratio $(n-i)/n$. Although nothing guarantees that the two caps do not overlap or, contrarily, leave blank a spherical strip in between them, at least the geometrical structure within each cap is preserved and the fitness of the child will only depend on the seam between the two caps. A similar recombination strategy is used in Ref. [20].
Following the same reasoning, we implement mutations by sorting a set by distance to the first vector, generating a random integer $i$ between $1$ and $n$ and a random phase $\theta\in[0,2\pi)$ and rotating the last $n-i$ vectors of the sorted set by the relative phase, i.e. $(z_{1},z_{2},\ldots,z_{d})\to(z_{1},z_{2}e^{i\theta},\ldots,z_{d}e^{i\theta})$. For $d=2$, this amounts to rotating the last $n-i$ vectors of the set an angle $\theta$ around the north pole.
@@ V.1 The Thomson problem
Figure 2: Results for the Thomson problem. a) Relative error between the energy in (17) for the frames obtained with the GA and the energies reported in Ref. [50]. b) Bloch-sphere representation of the MOF obtained with the GA for $n=24$, which forms a snub cube. Each blue dot represents an electron and the black lines connect nearest neighbors.
@@ V.2 General maximally orthogonal frames
It is important to realize that just because we associate certain properties to the minimizers of certain energy functions, this does not mean that the best way to numerically obtain a frame with a given property is to minimize the corresponding energy function. In particular, our testing on the GA has revealed that minimizing the coherence is not the best way to yield frames with low coherence. The minimization process tends to converge to local minima that are far from optimal. In contrast, minimizing a softer energy function, such as the Riesz $s$-energy or the $p$-frame potential with finite $s$ and $p$, which in principle offer no guarantee of leading to low-coherence (Grassmannian) frames, yields better results. With this, the theoretical one-to-one relationship between energy functions and properties gives way, in practice, to a situation where the choice of energy function is part of the tuning of the algorithm. The parameters $s$ or $p$ become hyperparameters of the GA, that can be tweaked in the search for frames that maximize certain properties, such as uniformity or coherence (See Sec. III.2 for a discussion of how this is also the case in SOTA algorithms for the search of optimal packings).
@@ V.2.2 Discussion of the results
Figure 3: Maximally orthogonal frames for $d=4$ and $s=8$. Each panel shows the overlaps among vectors of the best MOF out of ten runs of the GA for a certain size $n$. The horizontal colored lines indicate the applicable coherence bound. The panel title indicates whether it is a high-symmetry frame such as an ETF, maximal set of MUB or SIC-POVM. A box inside the panel indicates whether the frame is a $p$-design and the tolerance $\varepsilon$ with which the corresponding Welch bound (8) is saturated.
@@
It is interesting to observe the creation and destruction of certain structures as $n$ varies. After the set of MUB at $n=20$, particularly after $n=23$, the plot evidences that the vector space starts to be overcrowded with vectors. The frames from $n=24$ to $n=28$ exhibit very little structure in their overlaps. However, from $n=29$ we start to see a gradual sorting of the overlaps that culminates for $n=40$. Similarly, the transition from the SIC-POVM for $n=16$ to the set of MUB for $n=20$ is also interesting. When an additional vector is added from $n=16$ to $n=17$, $4$ vectors reposition to form one of the bases that will later constitute the set of MUB at $n=20$. A similar phenomenon seems to occur at $n=23$, but it does not manifest in the following values of $n$.
@@ Appendix B Visualizing uniformity: share of the vector space
Figure 7: Maximally orthogonal frames for $d=4$ and $s=8$. Each panel shows the normalized share of the vector space $\beta_{j}$ (21) of each vector the best MOF out of ten runs of the GA for a certain size $n$. The horizontal black line is a guide to the eye marking the value $1$. The panel title indicates whether it is a high-symmetry frame such as an ETF, maximal set of MUB or SIC-POVM. The standard deviation of the share $\sigma_{\beta}$ is indicated within each panel.
@@
Figure 8: Genetic algorithm stochastic results. Normalized share of the vector space $\beta_{j}$ (21) of the maximally orthogonal frames for $d=4$ and $s=8$ and $n=23,40,43$. Each column shows the result of a different run of the GA that optimizes a different metric. The horizontal black line is a guide to the eye marking the value $1$. The standard deviation of the share $\sigma_{\beta}$ is indicated within each panel.

### v2

@@ I Introduction
As orthogonality is a pairwise property, increasing the orthogonality between a given pair of vectors will often come at the expense of decreasing the orthogonality between other pairs of vectors. Depending on how much value one assigns to the orthogonality of each pair, one will favor one “maximally orthogonal” set of vectors or another. Thus, the first step in this paper is providing a formal definition for maximally orthogonal frames of a complex space of arbitrary dimension. Once armed with this definition, we discuss their properties and establish their relation with notable frame families such as tight and Grassmannian frames, equiangular tight frames [10], symmetric, informationally complete, positive operator-valued measurements (SIC-POVMs) [11], maximal sets of mutually unbiased bases [12, 13] and complex projective $t$-designs [14, 15]. We also show that finding maximally orthogonal frames constitutes, in most cases, a numerical optimization problem. We discuss the relation with established numerical problems such as the Thomson problem [16], the problem of finding optimal packings in complex projective space [17] and the problem of finding optimal reference states for quantum classifiers [18, 19]. Inspired by the use of a genetic algorithm for the Thomson problem [20], we present a genetic algorithm that is capable of generating maximally orthogonal frames of arbitrary size in a complex space of arbitrary dimension. Then, we visualize and measure the quality of the frames produced by the genetic algorithm. In the cases where the problem of finding maximally orthogonal frames is equivalent to the Thomson problem or the problem of finding optimal packings in complex projective space, we compare the genetic algorithm with state-of-the-art (SOTA) numerical methods.
@@ II.2.2 Riesz $s$-energy and asymptotic uniformity
The Riesz $s$-energy is invariant under changes in the global phase of the vectors, so we can identify equivalence classes for all unit vectors differing by a global phase, which essentially defines the complex projective space $\mathbb{C}P^{d-1}$. A parameterization of $\mathbb{C}P^{d-1}$ in real coordinates forms a manifold of dimension 2(d1)2(d-1). Following the poppy-seed bagel theorem, we will have to set $s\geq 2(d-1)$ to generate MOFs, $\bar{\Phi}_{d,n}[s]$, that are asymptotically uniform, in the sense of uniformly distributed on the manifold that parametrizes $\mathbb{C}P^{d-1}$, for $n\to\infty$ [23].
@@ II.2.4 $1$-designs are tight
The frame potential was introduced by Bennedeto and Fickus, originally with $p=1$, for the study of finite normalized tight frames [28]. They proved that ${\rm FP}_{1}$ has several convenient properties: its global minima are tight frames and saturate the Welch bound ${\rm FP}_{1}(\bar{\Phi}_{d,n}[p=1])=n^{2}/d-n$, and all its local minima are degenerate, i.e. they are global minima and thus tight frames and $1$-designs. This tells us that we can see tight frames as a particular instance of MOFs as defined in Proposition II.1. In fact, Benedetto and Fickus already refer to a set of vectors that minimize the frame potential ${\rm FP}_{1}$ as maximally orthogonal. Thus, they equate MOFs with tight frames. As argued above, it is natural to relax this equation and consider tight frames as just one case of many MOFs. Just like ${\rm FP}_{1}$ leads to MOFs that are tight, other weighting functions are equally valid and endow the corresponding MOFs with other useful properties, such as uniformity or low coherence. We mention in passing that another consequence of Benedetto and Fickus’s work is that finding tight frames is an easy task. A local optimization from an arbitrary initial condition suffices to construct tight frames numerically, although constructive analytical methods have also been devised [29, 30]. Other weighting functions do not lead to such an easily minimizable energy function, and thus obtaining the corresponding MOFs poses a numerical challenge.
@@ II.2.5 Coherence: the $p\to\infty$ and $s\to\infty$ limits
$\lim_{s\to\infty}({\rm RE}_{s})^{1/s}=m^{1/s}\max_{i\neq j}\frac{1}{D(|\phi_{i}\rangle,|\phi_{j}\rangle)}\,$, (9)
where $m$ is the multiplicity of the largest overlap. Therefore, only the maximum overlap(s) (minimum distance(s)) contribute to the energy and sets that only differ in submaximal overlaps will be degenerate. In this setting, the MOF will be given by
@@ II.2.6 Coherence and degeneracy
Figure 1: Degeneracy of the coherence. Bloch-sphere representation (top) and overlaps among vectors (bottom) of the different Grassmannian frames for $d=2$ and $n=5$. Column a) shows the distribution of vectors obtained by subtracting one vector from the equator from $\bar{\Phi}_{2,6}[s\to\infty]$. The frames depicted in columns b) and c) are obtained by continuously spreading the remaining three vectors at the equator until they form an equilateral triangle. Note that, for each $j$ in all three panels, there are 5 dots in the plot, each corresponding to the overlap of vector $j$ with every vector in the frame (including itself). Visually, there appears to be only three dots for each $j$ because there are only three distinct overlaps and so three of the dots coincide perfectly.
@@ II.2.7 Universal maximally orthogonal frames
The maximum number of MUB $\mathcal{M}(d)$ obeys $\mathcal{M}(d)\leq d+1$, with equality only if $d$ is an integer power of a primer number. Thus, if $d$ is an integer power of a prime it is possible to construct highly symmetric frames of $d(d+1)$ vectors, by combining the $d+1$ MUB, where every vector has only nearest-neighbors with overlap $1/\sqrt{d}$ (the vectors of the other bases) and next-nearest-neighbors with null overlap (the other vectors of the same basis). Despite not being ETFs, frames built from $d+1$ sets of MUB are Grassmannian and tight [17] and complex projective $2$-designs [27].
@@ IV The genetic algorithm
Recombination is implemented as a random crossover of vectors from two individuals. A trivial recombination could consist on generating a random integer $i$ between $1$ and $n$ to create a child with the $i$ first vectors of parent A and the last nin-i vectors form parent B. However, the vectors within a frame are, a priori, not sorted in any particular fashion, i.e. their order does not reflect any geometrical structure in the manifold that parametrizes the vectors. Therefore, the subset of vectors inherited from each parent would be random, consisting of vectors from parent A that could be arbitrarily close to vectors from parent B, and thus throw away most of the orthogonality gained over the course of the optimization process. Because of this, a trivial recombination like the one just described would, most likely, create a child with a fitness much lower than that of its parents. To prevent this, we sort the vectors of each parent by distance to the first vector of each set (which is fixed and the same for all sets). After the sort, the order of the vectors in a set does reflect some structural properties. Namely, the first vector of the set marks a reference point in the manifold that is common for all sets. After sorting, we apply the random recombination by drawing the first $i$ vectors from parent A, which are expected to be relatively closer to the reference point, and the last nin-i vectors from parent B, which are expected to be relatively far from the reference point. To better understand this process, let us consider the case $d=2$. Now the manifold is the Bloch sphere and the reference point is given by the vector $(1,0)$, which marks the north pole. Selecting the first $i$ vectors from parent A corresponds to selecting vectors covering a spherical cap around the north pole with an approximate height determined by the ratio $i/n$. Then, selecting the nin-i vectors from parent B corresponds to selecting vectors covering a spherical cap around the south pole with an approximate height determined by the ratio (ni)/n(n-i)/n. Although nothing guarantees that the two caps do not overlap or, contrarily, leave blank a spherical strip in between them, at least the geometrical structure within each cap is preserved and the fitness of the child will only depend on the seam between the two caps. A similar recombination strategy is used in Ref. [20].
Following the same reasoning, we implement mutations by sorting a set by distance to the first vector, generating a random integer $i$ between $1$ and $n$ and a random phase $\theta\in[0,2\pi)$ and rotating the last nin-i vectors of the sorted set by the relative phase, i.e. $(z_{1},z_{2},\ldots,z_{d})\to(z_{1},z_{2}e^{i\theta},\ldots,z_{d}e^{i\theta})$. For $d=2$, this amounts to rotating the last nin-i vectors of the set an angle $\theta$ around the north pole.
@@ V.1 The Thomson problem
Figure 2: Results for the Thomson problem. a) Relative difference between the energy (17) for the frames obtained with the GA and the energies reported in Ref. [50] up to $n=100$. b) For illustration purposes, the Bloch-sphere representation of the MOF obtained with the GA for $n=24$ is shown, which forms a snub cube. Each blue dot represents an electron and the black lines connect nearest neighbors.
@@ V.2 General maximally orthogonal frames
It is important to realize that just because we associate certain properties to the minimizers of certain energy functions, this does not mean that the best way to numerically obtain a frame with a given property is to minimize the corresponding energy function. In particular, our testing on the GA has revealed that minimizing the coherence is not the best way to yield frames with low coherence. The minimization process tends to converge to local minima that are far from optimal. As discussed in Sec. II.2.6, this is because the coherence disregards any submaximal overlaps between vectors. Accordingly, in the optimization process, there is no incentive to reposition vectors of a frame that are only involved in submaximal overlaps. This is true even if these vectors are far from their optimal position required to achieve minimal coherence. We see this phenomenon as akin to the vanishing gradient problem in neural networks. When the energy function becomes, or in this case is (by construction), insensitive to some of the optimization parameters, the optimization process is unable to tune these parameters, even if doing so sufficiently would eventually result in a lower energy and thus better performance. In contrast, minimizing a softer energy function, such as the Riesz $s$-energy or the $p$-frame potential with finite $s$ and $p$, which in principle offer no guarantee of leading to low-coherence (Grassmannian) frames, yields better results. With this, the theoretical one-to-one relationship between energy functions and properties gives way, in practice, to a situation where the choice of energy function is part of the tuning of the algorithm. The parameters $s$ or $p$ become hyperparameters of the GA, that can be tweaked in the search for frames that maximize certain properties, such as uniformity or coherence (See Sec. III.2 for a discussion of how this is also the case in SOTA algorithms for the search of optimal packings).
@@ V.2.2 Discussion of the results
Figure 3: Maximally orthogonal frames for $d=4$ and $s=8$. Each panel shows the overlaps among vectors of the best MOF out of ten runs of the GA for a certain size $n$. The horizontal colored lines indicate the applicable coherence bound. The panel title indicates whether it is a high-symmetry frame such as an ETF, a maximal set of MUB or a SIC-POVM. A box inside the panel indicates whether the frame is a $p$-design and the tolerance $\varepsilon$ with which the corresponding Welch bound (8) is saturated.
@@
It is interesting to observe the creation and destruction of certain structures as $n$ varies. After the set of MUB at $n=20$, particularly after $n=23$, the plot evidences that the vector space starts to be overcrowded with vectors. The frames from $n=24$ to $n=28$ exhibit very little structure in their overlaps. However, from $n=29$ we start to see a gradual sorting of the overlaps that culminates for $n=40$. Similarly, the transition from the SIC-POVM for $n=16$ to the set of MUB for $n=20$ is also interesting. When an additional vector is added from $n=16$ to $n=17$, 44 vectors reposition to form one of the bases that will later constitute the set of MUB at $n=20$. A similar phenomenon seems to occur at $n=23$, but it does not manifest in the following values of $n$.
@@ Appendix B Visualizing uniformity: share of the vector space
Figure 7: Maximally orthogonal frames for $d=4$ and $s=8$. Each panel shows the normalized share of the vector space $\beta_{j}$ (21) of each vector the best MOF out of ten runs of the GA for a certain size $n$. The horizontal black line is a guide to the eye marking the value 11. The panel title indicates whether it is a high-symmetry frame such as an ETF, maximal set of MUB or SIC-POVM. The standard deviation of the share $\sigma_{\beta}$ is indicated within each panel.
@@
Figure 8: Genetic algorithm stochastic results. Normalized share of the vector space $\beta_{j}$ (21) of the maximally orthogonal frames for $d=4$ and $s=8$ and $n=23,40,43$. Each column shows the result of a different run of the GA that optimizes a different metric. The horizontal black line is a guide to the eye marking the value 11. The standard deviation of the share $\sigma_{\beta}$ is indicated within each panel.